Baron Corporation has a target capital structure of 75 percent common stock, 10 percent preferred stock, and 15 percent debt. Its cost of equity is 11 percent, the cost of preferred stock is 5 percent, and the pretax cost of debt is 6 percent. The relevant tax rate is 23 percent. a. What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

aftertax cost of debt=6*(1-tax rate)

=6*(1-0.23)=4.62%

WACC=Respective costs*Respective weight

=(0.75*11)+(0.1*5)+(4.62*0.15)

=9.44%(Approx).